OpenStudy (volleyballlover55):

A waffle cone has a height of 7 inches and a diameter of 3 inches. What is the volume of ice cream that can be contained within the cone? Use 3.14 for pi. Round your answer to the nearest hundredth. 15.43 cubic inches 16.49 cubic inches 17.86 cubic inches 18.89 cubic inches

OpenStudy (volleyballlover55):

@johnweldon1993 @Mehek14

OpenStudy (erak):

Use the volume of a cone formula. \[V = 1/3 \pi r ^{2} \frac{ h }{ 3 }\]

OpenStudy (volleyballlover55):

@ShadowLegendX @rebeccaxhawaii